OpenStudy (anonymous):

4/x+1 - 4x/3x simplify Please

OpenStudy (anonymous):

is x+1 all in the denominator?

OpenStudy (anonymous):

yea so is 3x

myininaya (myininaya):

\[\frac{4}{x+1}-\frac{4}{3}*\frac{x+1}{x+1}\]

OpenStudy (anonymous):

the LCD is 3x(x+1) so this will become, (4(3x)-4x(x+1))/3x(x+1) 12x-4x^2-4x/3x(x+1) -4x^2+8x/3x(x+1) x(-4x+8)/3x(x+1) (-4x+8)/3(x+1)

myininaya (myininaya):

multiply first fraction by 3/3

myininaya (myininaya):

pat we don't need the x

myininaya (myininaya):

since x/x=1 assuming x isnt zero

OpenStudy (anonymous):

yeah, but i just included it to the LCD, do not worry, the answer will be the same.

myininaya (myininaya):

i know lol just more work

OpenStudy (anonymous):

thankyou!

OpenStudy (anonymous):

yeah, lol

OpenStudy (anonymous):

its for my algebra 2 exam